# TFusionOcc: Student’s t-Distribution Based Object-Centric Multi-Sensor Fusion Framework for 3D Occupancy Prediction

### v1

@@ IV-I1 Qualitative Results On SurroundOcc-nuScenes Dataset
IV-I2 Qualitative Results On nuScenes-C Dataset
Figure 14: Qualitative results compare against other SOTA algorithms under fog and snow scenarios, displayed in the top and bottom sections, respectively. Better viewed when zoomed in.
Since the nuScenes dataset does not contain fog and snow scenarios, which are common driving conditions in the real world, we adopted the nuScenes-C dataset to evaluate our model across three settings within the same scene under fog and snow conditions, with the severity of each condition set to hard, and compare the visualization results with other SOTA algorithms. The visualization results are shown in Figure 14, with the main area of the discrepancy highlighted in dark blue and dark red squares.
Under foggy conditions, the surround-view images suffer significant visibility degradation due to strong atmospheric scattering, leading to reduced illumination, low contrast, and severely blurred object boundaries, especially in distant regions of the scene. Compared to baseline SOTA algorithms, our method, especially the T-SQIW-25600 setting, successfully detects all dynamic pedestrian objects in the scene (as shown by the dark blue squares) and also accurately captures the background objects’ contour as well (as shown by the dark red squares), demonstrating the robustness of the proposed algorithm under foggy conditions. Nevertheless, in the snow condition, the snowfall captured in the surround-view image appears to introduce salt-and-pepper noise to the image, and snowflakes occlude some objects’ contours, even the main body parts. In this case, our proposed method is less tolerant of salt-and-pepper noise than the baseline SOTA algorithms.
IV-J Model Efficiency
@@
Method Modality Latency (ms) ($\downarrow$) Memory (GB) ($\downarrow$) Params SurroundOcc* [30] C 472 5.98 180.51M InverseMatrixVT3D* [17] C 447 4.41 67.18M OccFusion(C+R)* [18] C+R 588 5.56 92.71M OccFusion(C+L)* [18] C+L 591 5.56 92.71M OccFusion(C+L+R)* [18] C+L+R 601 5.78 114.97M OccCylindrical* [16] C+L 398 10.63 111.62M TFusionOcc (T-SQ-12800) C+L 278.46 3.68 95.62M TFusionOcc (T-SQ-IW-12800) C+L 281.47 3.88 95.96M TFusionOcc (T-P-25600) C+L 385.49 4.73 97.51M TFusionOcc (T-SQ-25600) C+L 388.05 4.13 97.59M TFusionOcc (T-SQ-IW-25600) C+L 391.06 4.04 98.24M
TABLE XII: Efficiency comparison of multi-modal 3D semantic occupancy prediction on the nuScenes validation set. Algorithm results marked with * were taken from the paper that introduced the model and were measured on a different GPU.
@@
We evaluated the real-time performance of our model in various settings on an RTX 4090 GPU and presented the results in Table XII. The experiment results show the trade-off between the total number of primitives and the model inference speed. More primitives require more rendering operations, leading to higher latency. Surprisingly, all five settings have similar total model parameters. Based on the main performance benchmark shown in Table II, the T-SQ-IW-12800 setting delivers the best balance between detection and real-time performance.
IV-K Ablation Study
IV-K1 Ablation Study On Skeleton Merge Module
@@
Main-Skele Augment Overlap Remove Range Filter IoU ($\uparrow$) mIoU ($\uparrow$) 46.24 (Baseline) 30.79 (Baseline) 46.20 (-0.04%) 30.70 (-0.09%) 46.07 (-0.17%) 30.76 (-0.03%) 46.08 (-0.16%) 30.69 (-0.1%) 41.96 (-4.28%) 29.92 (-0.87%) 29.35 (-16.89%) 22.31 (-8.48%)
@@
TABLE XIII: Ablation study on the Skeleton Merge Module. Main-Skele: Lidar-side main skeleton branch. Augment: Camera-side augment skeleton branch. $\uparrow$:the higher, the better.
@@
TABLE XIV: Ablation study on range filter hyperparameter $r$. $\uparrow$:the higher, the better.
We conducted an ablation study for each component of the skeleton merge module and demonstrated the results of the experiment in Table XIII. The experiment results prove the significant impact of the main skeleton structure, which supports the fundamental 29.92% mIoU and 41.96% IoU performance, and the auxiliary impact of the augment structure, which further boosts 0.87% mIoU and 4.28% IoU performance. The main skeleton, based on lidar, provides detailed 3D geometric information of the surrounding scene, and the augmented skeleton, based on the camera, provides additional information to mitigate occlusion impact in the scene. Furthermore, we studied the impact of the range filter hyperparameter $r$ on the overall performance of the model and present the results in Table XIV. The results show that a smaller filter range forces the augmented skeleton to get too close to the main skeleton, preventing it from overcoming the lidar sensor’s occlusion limitation.
IV-K2 Ablation Study On Depth-Fusion-Based 3D Deformable Attention Module
@@
TABLE XV: Ablation study on Depth-Fusion-Based 3D Deformable Attention Module. $\uparrow$:the higher, the better. Vis-Depth Fusion: add the fused depth feature back to the visual feature, resulting in a depth-aware visual feature. Depth-Fusion: fuse the visual depth feature with the lidar depth feature. Vis-Depth: visual depth feature. Lidar-Depth: lidar depth feature. $\uparrow$:the higher, the better.
@@
We conducted an ablation study with respect to each component of the depth-fusion-based 3D deformable attention module and presented the experiment results in Table XV. In practice, Vis-Depth provides dense but less accurate depth maps, while Lidar-Depth provides sparse but more accurate depth maps. The experiment shows that dense depth maps contribute significantly to the performance, even though they are less accurate. Relying solely on sparse but accurate depth maps from lidar results in a significant performance degradation of 5.5%-6.0%. Furthermore, adding fused depth maps back to the original visual features results in depth-aware visual features contributing the second-largest performance improvement, around 2.2%. The experiment results demonstrate the necessity of multi-stage multi-modality feature fusion to boost model performance.
IV-K3 Ablation Study On MGCAFusion Module
@@
Skip-Connect WSFusion GCFusion IoU ($\uparrow$) mIoU ($\uparrow$) 46.24 (Baseline) 30.79 (Baseline) 38.60 (-7.64%) 18.18 (-12.61%) 39.39 (-6.85%) 18.57 (-12.22%) 24.20 (-22.04%) 19.91 (-10.88%)
@@
TABLE XVI: Ablation study on the MGCAFusion Module. Skip-Connet: skip-connection structure at the last feature fusion stage. WSFusion: weighted summation fusion part. GCFusion: gated concatenation fusion part. $\uparrow$:the higher, the better.
@@
We conducted an ablation study for each major component of the MGCAFusion module and present the experiment results in Table XVI. The experiment results show that the weighted summation and gated concatenation fusion components contribute almost equally to the model’s final performance, with the former slightly more important. Missing each part will cause 12.22% to 12.61% mIoU and 6.85% to 7.64% IoU performance degradation. The skip-connection structure plays an important role during the feature fusion.
V Conclusion and Future work
This paper presents TFusionOcc, a novel Student’s t-distribution-based object-centric multi-stage multi-sensor fusion framework. Our framework in the sensor fusion part adopts a multi-stage feature-level fusion strategy, including an early-stage, a middle-stage, and a late-stage fusion, to better integrate complementary information across modalities and mitigate the drawbacks of each modality. We also adopt an object-centric modeling strategy by leveraging the Student’s t-distribution and the T-Mixture model as a probability kernel, combined with several typical primitives, such as the general T-Primitive and the superquadric, to achieve greater robustness against outliers. We further propose a primitive called the deformable superquadric primitive (superquadric with inverse-warp), which offers greater geometric shape representation flexibility, thereby boosting the model’s 3D geometric shape-capturing capability in the driving scene. The extensive experiments conducted on the nuScenes [2] and nuScenes-C [31] datasets, including challenging cases of rainy, night, fog, snow and sensor-malfunction, demonstrate that our proposed approach not only achieves SOTA performance but also exhibits superior robustness. In our future research endeavors, we will investigate how to utilize a much smaller set of primitives and fewer rendering operations to achieve similar performance.

### v2

@@ IV-G Qualitative Study
In the daytime scenario, as shown at the top of Fig. 10, in the orange square highlighted area, benefit from multi-task learning, which provides an additional supervision signal, DAOcc successfully detects distant pedestrians and achieves similar performance as TFusionOcc. However, in the dark-red-square-highlighted area, DAOcc fails to detect the distant vehicle, demonstrating the superior performance of our proposed method. In the rainy scenario, as shown in the middle of Fig. 10, in the purple square highlighted area, our proposed method not only successfully detects all dynamic objects, but also preserves their general contours well. In the highlighted dark-blue-square area, our approach achieves par performance with other SOTA algorithms for detecting challenging pedestrians. In the nighttime scenario, as shown at the bottom of Fig. 10, our proposed method successfully detects very distant dynamic objects in the highlighted dark-red-square area, demonstrating its superior long-range dynamic object perception capability.
IV-H Model Efficiency
@@
Method Modality Latency (ms) ($\downarrow$) Memory (GB) ($\downarrow$) Params M-CONet [30] C+L 416.00 12.10 125.43M OccCylindrical [16] C+L 361.39 5.97 111.62M OccMamba-384 [10] C+L 448.30 12.1 111.08M OccMamba-128 [10] C+L 269.30 6.8 64.58M TFusionOcc (T-SQ-12800) C+L 278.46 3.68 95.62M TFusionOcc (T-SQ-IW-12800) C+L 281.47 3.88 95.96M TFusionOcc (T-P-25600) C+L 385.49 4.73 97.51M TFusionOcc (T-SQ-25600) C+L 388.05 4.13 97.59M TFusionOcc (T-SQ-IW-25600) C+L 391.06 4.04 98.24M
TABLE VI: Model efficiency comparison of different methods. The experiments are performed on a single RTX4090 using six multi-camera images and 10 sweeps of LiDAR. For input image resolution, all methods adopt $1600\times 900$.
@@
We evaluated the real-time performance of our model across various settings, comparing it with several other SOTA approaches on an RTX 4090 GPU, and presented the results in Table VI. The experiment results show the trade-off among the total number of primitives, memory consumption, and model inference speed. More primitives require more rendering operations and memory, leading to higher latency but better performance. Based on the main performance benchmark shown in Table II, the T-SQ-IW-12800 setting delivers the best balance between prediction and real-time performance.
IV-I Ablation Study
IV-I1 Ablation Study on Different Distribution and Primitives
Prob Distribution Query Number IoU ($\uparrow$) Gain ($\uparrow$) mIoU ($\uparrow$) Gain ($\uparrow$) T-P / G-P 25600 44.77 / 44.71 – / – 29.22 / 29.20 – / – T-SQ / G-SQ 25600 45.40 / 44.51 +0.63 / -0.20 30.38 / 30.25 +1.16 / +1.05 T-SQ-IW / G-SQ-IW 25600 46.23 / 45.58 +0.83 / +1.07 30.77 / 30.66 +0.39 / +0.41
TABLE VII: Ablation study on different probability distribution and primitives. Gain denotes the vertical performance change within each series (T-series / G-series). T refers to the Student’s t-distribution kernel, and G refers to the Gaussian kernel.
Table VII presents the ablation study results for different probability distributions and primitive settings. For each row in the table, for the same type of primitive and the same total number of primitives, with different distribution settings, the model obtained better performance, benefiting from the robustness property of the Student’s t-distribution. Notably, the geometric prediction capability benefits more than the semantic prediction capability. What’s more, by examining the table vertically, we can see that, regardless of the probability distribution kernel setting, the primitive with greater shape variety performs better, demonstrating the importance of the primitive’s geometric shape representation capability.
IV-I2 Ablation Study on Basis Field of T-SQ-IW Primitive
Basis field IoU ($\uparrow$) mIoU ($\uparrow$) $B_{1}\sim B_{24}$ 46.23 (Baseline) 30.77 (Baseline) $B_{10}\sim B_{24}$ 32.98 (-13.25%) 23.17 (-7.60%) $B_{1}\sim B_{18}$ 36.20 (-10.03%) 24.11 (-6.66%) $B_{1}\sim B_{9}$ 21.83 (-24.40%) 11.60 (-19.17%) No Inverse-Warp 20.24 (-25.99%) 13.87 (-16.90%)
TABLE VIII: Ablation study on basis field in T-SQ-IW primitive. $B_{10}\sim B_{24}$: primitive loses basic constant, linear, shear and taper deformation capability. $B_{1}\sim B_{18}$: primitive loses quadratic, radial bulge and smooth corner deformation capability. $B_{1}\sim B_{9}$: primitive further loses shear, twist and bending deformation capability.
Table VIII presents the results of the ablation study for each set of basis fields in the primitive T-SQ-IW. Since $B_{19}\sim B_{24}$ provides the highest level deformation, including quadratics, radial bulge, and smooth corner deformation, losing those basis fields causes 10.04% IoU and 6.66% mIoU degradation. In addition, because $B_{10}\sim B_{18}$ provides the second-highest level of deformation, including full shear, twist, and bending, losing $B_{10}\sim B_{24}$ results in performance drop. Surprisingly, the performance drop after losing all basis fields is similar to the loss $B_{10}\sim B_{24}$. This result is very likely because the remaining $B_{1}\sim B_{9}$ only provide basic constant, linear, shears, and taper deformation, which leads to limited improvement in the shape variety of the primitive. This result also demonstrates the importance of the high-level deformation basis field, which significantly improved the shape variety of the primitive, resulting in better overall 3D occupancy prediction performance.
IV-I3 Ablation Study on Skeleton Merge Module
@@
Main-Skele Augment Overlap Remove Range Filter IoU ($\uparrow$) mIoU ($\uparrow$) 46.24 (Baseline) 30.79 (Baseline) 46.20 (-0.04%) 30.70 (-0.09%) 46.07 (-0.17%) 30.76 (-0.03%) 46.08 (-0.16%) 30.69 (-0.10%) 41.96 (-4.28%) 29.92 (-0.87%) 29.35 (-16.89%) 22.31 (-8.48%)
@@
TABLE IX: Ablation study on the Skeleton Merge Module. Main-Skele: LiDAR-side main skeleton branch. Augment: Camera-side augment skeleton branch.
@@
TABLE X: Ablation study on range filter hyperparameter $r$.
We conducted an ablation study for each component of the skeleton merge module and presented the results in Table IX. The experiment results prove the significant impact of the main skeleton structure, which supports the fundamental 29.92% mIoU and 41.96% IoU performance, and the auxiliary impact of the augmented skeleton, which further boosts 0.87% mIoU and 4.28% IoU performance. The main skeleton, based on LiDAR, provides precise 3D geometric information about the surrounding scene, especially the front faces of objects, and the augmented skeleton, based on the camera, provides additional information to mitigate the partial occlusion. Furthermore, we studied the impact of the range filter hyperparameter $r$ on the model’s overall performance and present the results in Table X. The results show that a smaller filter range forces the augmented skeleton to get too close to the main skeleton, preventing it from overcoming the LiDAR sensor’s occlusion limitation, resulting in performance degradation.
IV-I4 Ablation Study on Fused Depth Map Guided 3D Deformable Attention Module
@@
TABLE XI: Ablation study on Depth-Fusion-Based 3D Deformable Attention Module. Vis-Depth Fusion: add the fused depth maps back to the visual features, resulting in depth-aware visual features. Depth-Fusion: fuse the visual depth maps with the LiDAR depth maps. Vis-Depth: visual depth maps. LiDAR-Depth: LiDAR depth maps.
@@
We conducted an ablation study for each component of the fused-depth-guided 3D deformable attention module and presented the results in Table XI. In practice, Vis-Depth provides dense but less accurate depth maps that are good at capturing the partial occlusion of object structure. In contrast, LiDAR-Depth provides sparse but more accurate depth maps, especially good at capturing the front faces of objects in the scene. The experiment shows that Vis-Depth maps contribute significantly to the performance, even though they are less accurate. Relying solely on sparse but accurate LiDAR-Depth maps results in a performance degradation of 6.44% IoU and 5.51% mIoU, respectively. Furthermore, adding fused depth maps back to the original visual features results in depth-aware visual features contributing the second-largest performance improvement. The experiment results demonstrate the importance of depth map density and the necessity of multi-source depth map fusion and vis-depth feature fusion to boost model performance.
IV-I5 Ablation Study on GCWSFusion Module
We conducted an ablation study for each major component of the GCWSFusion module and present the experiment results in Table XII.
@@ IV-I5 Ablation Study on GCWSFusion Module
Gate-atten Weight-sum Skip-Connect WSFusion GCFusion IoU ($\uparrow$) mIoU ($\uparrow$) 46.24 (Baseline) 30.79 (Baseline) 45.52 (-0.72%) 30.68 (-0.11%) 41.34 (-4.90%) 27.12 (-3.67%) 38.60 (-7.64%) 18.18 (-12.61%) 39.39 (-6.85%) 18.57 (-12.22%) 24.20 (-22.04%) 19.91 (-10.88%)
@@
TABLE XII: Ablation study on the GCWSFusion Module. Skip-Connet: skip-connection structure at the last feature fusion stage. WSFusion: weighted summation fusion part. GCFusion: gated concatenation fusion part. Gate-atten: Gate attention mechanism in GCFusion module. Weight-sum: Weighted summation mechanism in the WSFusion module.
@@
The results of the experiment show that the weight-summation mechanism plays a relatively more important role than the gate-attention mechanism in the overall GCWSFusion module. However, the WSFusion and GCFusion components contribute almost equally to the model’s final performance, with WSFusion slightly more important. Missing each part will cause 12.22% to 12.61% mIoU and 6.85% to 7.64% IoU performance degradation. The skip-connection structure also plays an important role during feature fusion; without it, the mIoU and IoU degrade by 22.04% and 10.88%, respectively.
V Conclusion
This paper presents TFusionOcc, a T-primitive-based object-centric multi-sensor fusion framework for 3D semantic occupancy prediction. The proposed method is built upon a family of Student’s t-distribution-based T-primitives, among which the deformable T-Superquadric with inverse warping serves as the key geometry-enhancing primitive. To further improve robustness, we introduce a unified probabilistic formulation based on Student’s t-distribution and the T-mixture model for joint occupancy and semantic modeling, together with a tightly coupled multi-stage fusion architecture for effective camera–LiDAR integration. Extensive experiments on SurrounOcc-nuScenes, Occ3D-nuScenes, and nuScenes-C demonstrate that TFusionOcc achieves strong performance and robustness in the most challenging conditions. In future work, we plan to further improve long-range perception and robustness under structured visual degradations, such as motion blur and snow.